A coin bank contains 25 dimes and nickels. The total amount of money in the bank is $2.15. Find the number of dimes in the bank.
Answer by Boreal(15235)

x dimes
25-x nickels
.10x+.05(25-x)=2.15
.10x-.05x+1.25=2.15
.05x=0.90
x=18 dimes $1.80
25-x=7 nickels $0.35. They add to $2.15